OpenStudy (anonymous):

determine the slope of the line given the following equations how to apply it here ??? 2x=3y+7???

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq.htm

OpenStudy (anonymous):

so its -3y=-2x+7?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

after that ??

OpenStudy (anonymous):

wait let me call some friends

OpenStudy (anonymous):

k

OpenStudy (anonymous):

put the equation in this form y=mx+b can you do that ?!

OpenStudy (anonymous):

-3y=-2x+7

OpenStudy (anonymous):

you have -3y...how to turn it y=...

OpenStudy (anonymous):

multiply by -3?

OpenStudy (anonymous):

No :'( ! by \[-\frac{1}{3}\]

OpenStudy (anonymous):

why??

OpenStudy (anonymous):

Good question ! we had -3y \[-\frac{1}{3}(-3y)\]=?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

=y

OpenStudy (anonymous):

yes ! "y" that's what we are looking for ! multiply by \[-\frac{1}{3}\] both sides of the equation !

OpenStudy (anonymous):

so y=y+7/3?

OpenStudy (anonymous):

see now this form" y=(slope)x+b" ?!

OpenStudy (anonymous):

no ! you had "-3y=-2x+7"

OpenStudy (anonymous):

y=-2x+7

OpenStudy (anonymous):

multiply by \[-\frac{1}{3}\] both sides ! and you'll get y=(slope)x+b Sorry I have to go ! you just multiplied one side ! y=\[-\frac{1}{3} (-3y) =(\frac{-1}{3})(-2x+7)\]=(slope)x+b

OpenStudy (anonymous):

-2/3x+-7/3?

OpenStudy (anonymous):

+2/3x-7/3 ...negative by negative is positive @alloids65432